# Causal Representation Learning from Multimodal Biomedical Observations

### v2

@@ 5 Estimation Model Architectures
We use flow to estimate the exogenous variables $\bm{\epsilon}$ in Eq. (1) and implement the causal relations through a learnable adjacency matrix $\hat{\mathbf{A}}$. The binary values in $\hat{\mathbf{A}}$ represent the causal generation process between latent variables, e.g. $\hat{A}_{i,j}=1$ indicates $\hat{z}_{j}$ is the parent of $\hat{z}_{i}$, while $\hat{A}_{i,j}=0$ means $\hat{z}_{j}$ dose not contribute to the generation of $\hat{z}_{i}$. For each component $\hat{z}_{i}$, we select its parents $\text{Pa}(\hat{z}_{i})$ based on the adjacency matrix, and apply the flow transformation to $\hat{\epsilon}_{i}$.
@@ 6.1 Numerical Dataset
In the numerical simulations, we consider three cases with different numbers of modalities and inter-modal causal relations. Case 1: 15-dimensional observations across two modalities, each with two latent and one exogenous variable. Case 2: 20-dimensional observations across two modalities, each with three latent and one exogenous variable. Case 3: 15-dimensional observations across four modalities, each with two latent and one exogenous variable. The nonparametric mixing function is simulated by a random MLP with LeakyReLU units, and the inter-modal latent variables are sparse causally related. The detailed data-generation process is provided in Appendix D.1.
@@
As an ablation study, we further show the consequences of violating the sparsity assumptions to validate our theorem. Based on case 1, we create four types of datasets with different sparsity ratios and report the MCC in each scenario in Figure 5(c). The sparsity ratio represents the ratio of existing causal links to all possible causal links between modality-specific latent variables. A value of 0 indicates that the latent variables between modalities are fully connected, while higher values correspond to sparser connections. The result shows that identifiability can be better achieved with a higher sparsity ratio, and our framework outperforms other baselines in all scenarios.
@@ D.1 Numerical Dataset
We generate $n=10000$ samples according to Eq. (1) and Eq. (2). The dimensionality of the observations in each modality ranges from $d(\mathbf{x})=15$ to $20$. Following prior work (Von Kügelgen et al., 2021; Yao et al., 2021; Zimmermann et al., 2021), we generate observations using a multi-layer perceptron (MLP). Specifically, the mixing function $g$ is modeled as a three-layer MLP with randomly initialized weights and leaky ReLU activations, enabling $g$ to represent a general nonparametric mixing function. The causal noise terms $\epsilon$ are independently and identically distributed (i.i.d.), and the exogenous variables are mutually independent. Sparse inter-modality causal dependencies are randomly generated, ensuring that each modality’s latent variables maintain at least one causal connection with another modality.
@@
We generate two 15-dimensional modality observations under different sparsity ratios. Each observation is generated from three latent variables: two causally related latent variables and one exogenous variable. The sample size for each dataset is set to $n=10000$, and the dimensionality of the observations in each modality is $d(\mathbf{x})=15$. Let the four latent variables be denoted as $z^{(1)}_{1},z^{(1)}_{2},z^{(2)}_{1}$, and $z^{(2)}_{2}$. The sparsity ratio determines the extent of inter-modality connections among these latent variables: For sparsity ratio = 0%, all inter-modality latent variables are fully connected. For sparsity ratio = 25%, three pairs of inter-modality connections among the latent variables exist. For sparsity ratio = 50%, two pairs of inter-modality latent variables are connected. For sparsity ratio = 75%, only one inter-modality latent pair remains connected.
@@ Appendix E Extended Experiment
To evaluate the scalability and generalizability of our method to complex causal structures, we conducted additional experiments on higher-dimensional simulated tasks with diverse configurations of latent variables and modalities. These setups introduce significantly more complex causal relationships between variables. Specifically, we consider three extended scenarios: (1) Two-mods, involving 30-dimensional observations from two modalities with four latent variables and one exogenous variable per modality. (2) Five-mods, with 30-dimensional observations from five modalities with two latent variables and one exogenous variable per modality. (3) Six-mods, with 30-dimensional observations from six modalities with two latent variables and one exogenous variable per modality. The results, summarized in Table 4, show that our method consistently delivers robust performance under these challenging conditions.
@@
Two mods Five mods Six mods Cyclic within Cyclic across Two mods Three mods
R2 0.89$\pm$ 1e-4 0.89$\pm$ 1e-4 0.97$\pm$ 8e-7 0.95$\pm$ 1e-5 0.94$\pm$ 2e-4 0.86$\pm$ 5e-4 0.90$\pm$ 1e-4
MCC 0.83$\pm$ 4e-6 0.84$\pm$ 3e-4 0.82$\pm$$\pm$ 4e-4 0.89$\pm$ 2e-4 0.92$\pm$ 1e-5 0.83$\pm$$\pm$ 7e-4 0.83$\pm$ 4e-6
@@ F.2 Training Details
Hyperparameter Details.
The hyperparameters $\alpha=[\alpha_{\text{Ind}},\alpha_{\text{Sp}},\alpha_{\text{Recon}}]$ represent the weights assigned to each term in the composite objective function. For the experiments, the following settings were applied: $\alpha=[1e-2,1,10]$ for the synthetic dataset, $\alpha=[1e-5,2,10]$ for the MNIST dataset, and $\alpha=[1e-1,1e-2,1]$ for the phenotype dataset.
@@ Appendix G Algorithm Pseudocode
20:  Choose the parents of each latent variable $\hat{z}_{i}$ based on the adjacency matrix
21:  for each flow block do
22:     Pass $\text{Pa}(\hat{z}_{i})$ through flow to obtain estimated residuals $\hat{\epsilon}_{i}$
23:     Update the estimated causal graph based on the adjacency matrix with the threshold
24:     Compute sparsity loss based on $L_{1}$ norm
25:  end for
26:  Compute the KL divergence between [$[\{\hat{\eta}^{(m)}\}_{m=1}^{M},\hat{\epsilon}^{d(\mathbf{z})}_{i=1}]$ and Gaussian prior
27:  return Estimated noise term $\hat{\epsilon}^{d(\mathbf{z})}_{i=1}$
28:
29:  # Decoder
30:  Input: Estimated latent and exogenous variables in each group $\{\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)}\}_{m=1}^{M}$
31:  Output: Reconstructed grouped features $\{\hat{\mathbf{x}}^{(m)}\}_{m=1}^{M}$
32:  for each group $m=1$ to $M$ do
33:     Decode $(\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)})$ to reconstruct features $\hat{\mathbf{x}}^{(m)}$: $\hat{\mathbf{x}}^{(m)}=\text{De}^{(m)}(\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)})$
34:     Compute reconstruction loss using MSE: $\mathcal{L}_{\text{Recon}}^{(m)}=\text{MSE}(\hat{\mathbf{x}}^{(m)},\mathbf{x}^{(m)})$
35:  end for

### v3

@@ 5 Estimation Model Architectures
We use flow to estimate the exogenous variables $\bm{\epsilon}$ in Eq. (1) and implement the causal relations through a learnable adjacency matrix $\hat{\mathbf{A}}$. The binary values in $\hat{\mathbf{A}}$ represent the causal generation process between latent variables, e.g. $\hat{A}_{i,j}=1$ indicates $\hat{z}_{j}$ is the parent of $\hat{z}_{i}$, while $\hat{A}_{i,j}=0$ means $\hat{z}_{j}$ dose not contribute to the generation of $\hat{z}_{i}$. For each component $\hat{z}_{i}$, we select its parents $\text{Pa}(\hat{z}_{i})$ based on the adjacency matrix, and apply the flow transformation to get $\hat{\epsilon}_{i}$.
@@ 6.1 Numerical Dataset
In the numerical simulations, we consider three cases with different numbers of modalities and inter-modal causal relations. Case 1: 15-dimensional observations across two modalities, each with two latent and one exogenous variable. Case 2: 20-dimensional observations across two modalities, each with three latent and one exogenous variable. Case 3: 30-dimensional observations across four modalities, each with two latent and one exogenous variable. The nonparametric mixing function is simulated by a random MLP with LeakyReLU units, and the inter-modal latent variables are sparse causally related. The detailed data-generation process is provided in Appendix D.1.
@@
As an ablation study, we further show the consequences of violating the sparsity assumptions to validate our theorem. Based on case 2, we create four types of datasets with different sparsity ratios and report the MCC in each scenario in Figure 5(c). The sparsity ratio represents the ratio of existing causal links to all possible causal links between modality-specific latent variables. A value of 0 indicates that the latent variables between modalities are fully connected, while higher values correspond to sparser connections. The result shows that identifiability can be better achieved with a higher sparsity ratio, and our framework outperforms other baselines in all scenarios.
@@ D.1 Numerical Dataset
We generate $n=10000$ samples according to Eq. (1) and Eq. (2). Following prior work (Von Kügelgen et al., 2021; Yao et al., 2021; Zimmermann et al., 2021), we generate observations using a multi-layer perceptron (MLP). Specifically, the mixing function $g$ is modeled as a three-layer MLP with randomly initialized weights and leaky ReLU activations, enabling $g$ to represent a general nonparametric mixing function. The causal noise terms $\epsilon$ are independently and identically distributed (i.i.d.), and the exogenous variables are mutually independent. Sparse inter-modality causal dependencies are randomly generated, ensuring that each modality’s latent variables maintain at least one causal connection with another modality.
@@
For the ablation study, we generate two modality observations under different sparsity ratios. Each observation is generated from three causally related latent variables and one exogenous variable. The sample size for each dataset is set to $n=10000$, and the dimensionality of the observations in each modality is $d(\mathbf{x})=20$. The sparsity ratio determines the extent of inter-modality connections among these latent variables. A higher sparsity ratio leads to a sparser causal structure, meaning fewer causal connections between latent variables. Conversely, a lower sparsity ratio yields a denser causal matrix with more causal dependencies. For example, a sparsity ratio of 0% indicates that all inter-modality latent variables are fully connected, whereas a sparsity ratio of 50% implies that half of the possible causal edges are removed.
@@ Appendix E Extended Experiment
To evaluate the scalability and generalizability of our method to complex causal structures, we conducted additional experiments on higher-dimensional simulated tasks with diverse configurations of latent variables and modalities. These setups introduce significantly more complex causal relationships between variables. Specifically, we consider three extended scenarios: (1) Five-mods, with 30-dimensional observations from five modalities with two latent variables and one exogenous variable per modality. (2) Six-mods, with 30-dimensional observations from six modalities with two latent variables and one exogenous variable per modality. (3) Eight-mods, involving 30-dimensional observations from eight modalities with two latent variables and one exogenous variable per modality. The results, summarized in Table 4, show that our method consistently delivers robust performance under these challenging conditions.
@@
Five mods Six mods Eight mods Cyclic within Cyclic across Two mods Three mods
R2 0.89$\pm$ 1e-4 0.97$\pm$ 8e-7 0.83$\pm$ 1e-3 0.95$\pm$ 1e-5 0.94$\pm$ 2e-4 0.86$\pm$ 5e-4 0.90$\pm$ 1e-4
MCC 0.84$\pm$ 3e-4 0.82$\pm$ 4e-4 0.91$\pm$ 5e-4 0.89$\pm$ 2e-4 0.92$\pm$ 1e-5 0.83$\pm$$\pm$ 7e-4 0.83$\pm$ 4e-6
@@ F.2 Training Details
Hyperparameter.
The hyperparameters $\alpha=[\alpha_{\text{Ind}},\alpha_{\text{Sp}},\alpha_{\text{Recon}}]$ represent the weights assigned to each term in the composite objective function. For each dataset, they were tuned within appropriate logarithmic intervals, ensuring a balance between independence, sparsity, and reconstruction. For the experiments, the following settings were applied: $\alpha=[1e-1,1e-2,1]$ for the synthetic dataset, $\alpha=[1e-2,1e-3,2]$ for the MNIST dataset, and $\alpha=[1e-1,1e-2,1]$ for the phenotype dataset.
@@ Appendix G Algorithm Pseudocode
20:  Select the parents of latent variable based on the adjacency matrix
21:  Pass through flow model to obtain estimated residuals $\hat{\epsilon}_{i}$
22:  Update the estimated causal graph based on the adjacency matrix with threshold
23:  Compute sparsity loss based on $L_{1}$ norm
24:  Compute the KL divergence between [$[\{\hat{\eta}^{(m)}\}_{m=1}^{M},\hat{\epsilon}^{d(\mathbf{z})}_{i=1}]$ and Gaussian prior
25:  return Estimated noise term $\hat{\epsilon}^{d(\mathbf{z})}_{i=1}$
26:
27:  # Decoder
28:  Input: Estimated latent and exogenous variables in each group $\{\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)}\}_{m=1}^{M}$
29:  Output: Reconstructed grouped features $\{\hat{\mathbf{x}}^{(m)}\}_{m=1}^{M}$
30:  for each group $m=1$ to $M$ do
31:     Decode $(\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)})$ to reconstruct features $\hat{\mathbf{x}}^{(m)}$: $\hat{\mathbf{x}}^{(m)}=\text{De}^{(m)}(\hat{\mathbf{z}}^{(m)},\hat{\eta}^{(m)})$
32:     Compute reconstruction loss using MSE: $\mathcal{L}_{\text{Recon}}^{(m)}=\text{MSE}(\hat{\mathbf{x}}^{(m)},\mathbf{x}^{(m)})$
33:  end for